You have the following initial information on CMR Co. on which to base your calculations and discussion for questions 1) and 2):
• Current long-term and target debt-equity ratio (D:E) = 1:4
• Corporate tax rate (TC) = 30%
• Expected Inflation = 1.75%
• Equity beta (E) = 1.6385
• Debt beta (D) = 0.2055
• Expected market premium (rM – rF) = 6.00%
• Risk-free rate (rF) = 2.15%


1) The CEO of CMR Co., for which you are CFO, has requested that you evaluate a potential investment in a new project. The proposed project requires an initial outlay of $7.15 billion. Once completed (1 year from initial outlay) it will provide a real net cash flow of $575 million in perpetuity following its completion. It has the same business risk as CMR Co.’s existing activities and will be funded using the firm’s current target D:E ratio.
a) What is the nominal weighted-average cost of capital (WACC) for this project? (3 marks)
b) As CFO, do you recommend investment in this project? Justify your answer (numerically). (2 marks)
Answer #1
  1. Cost of Equity = (1.6385 X 6%) + 2.15% = 11.98% ~ 12%
  2. Assume cost of debt to be 3.4% {(0.2055 X 6%) + 2.15%}

Cost after tax benefit = 1 - 0.30 = 0.70. Therefore = 3.4% X 0.70 = 2.38%

Now,

WACC = 0.75 X 11.98% +  0.25 X 2.38%

= 8.98% + 0.6%

= 9.58%

NPV for project = $575M / 1.75% X 100 = $32.85 B

Hence, take the project!